Suppose the demand function that firm faces shifted from Qd=100-2p to Qd=80-3p and the supply function has shifted fromQs=50+4p to Qs=80+5pso that find the effect of this change on priceand quantity? and Which of the changes in demand and supply is higher?

1
Expert's answer
2021-12-02T14:08:30-0500

At equilibrium; "Q_{d}=Q_{s}"


"100-2p=50+4p"


"50=6p"


"p=8.33"

Change in quantity demanded;

"=(80-3p)-(100-2p)"

"=(80-3(8.33))-(100-2(8.33))=-28.33"


Change in quantity supplied;

"=(80+5p)-(50+4p)"

"=(80+5(8.33))-(50+4(8.33))=104.97"


The change in supply is higher